### v1

# Connecting Discrete Morse Functions via Birth-Death Transitions

Abstract.
We study transformations between discrete Morse functions on a finite simplicial complex via birth-death transitions—elementary chain maps between discrete Morse complexes that either create or cancel pairs of critical simplices. We prove that any two discrete Morse functions $f_{1},f_{2}:K\to\mathbb{R}$ on a finite simplicial complex $K$ are linked by a finite sequence of such transitions. As applications, we present alternative proofs of several of Forman’s fundamental results in discrete Morse theory and study the topology of the space of discrete Morse functions.
Key words and phrases:
Discrete Morse theory, birth-death transitions, simplicial complexes, complex of discrete Morse functions.
1. Introduction
In classical Morse theory, a well-known result states that two Morse functions $\phi_{1},\phi_{2}:M\to\mathbb{R}$ on a compact manifold $M$ can be continuously deformed into one another via a one-parameter family [3, 2]. This fundamental result plays a crucial role in understanding the structure of differentiable manifolds by providing a mechanism to relate different Morse functions on the same manifold.
In discrete Morse theory, a parallel result is naturally expected. Discrete Morse theory, introduced by Forman [7, 8], adapts the philosophy of smooth cases to simplicial complexes and cell complexes, replacing smooth functions with combinatorial analogs. For a given simplicial complex $K$, a discrete Morse function assigns values to its simplices in a manner compatible with topological gradient flow.
Let $K$ be a simplicial complex and $f_{1},f_{2}:K\to\mathbb{R}$ be two discrete Morse functions. In [10] and our previous works [16, 17], we introduced the concept of connectedness between both critical simplices and discrete Morse functions enabling an algebraic characterization of the relationship between their associated gradient vector fields. In particular, the connectedness homomorphisms between discrete Morse complexes $C_{\ast}^{f_{1}},C_{\ast}^{f_{2}}$, corresponding to $f_{1}$ and $f_{2}$ respectively, provide a powerful tool for analyzing the relationships between $f_{1}$ and $f_{2}$.
Another central theme in discrete Morse theory is the study of the complex of discrete Morse functions, denoted $\mathcal{M}(K)$, as introduced by Chari and Joswig [4]. The authors constructed $\mathcal{M}(K)$ as a simplicial complex to encode the space of discrete Morse functions on $K$, establishing a bijection between $\mathcal{M}(K)$ and the set of gradient vector fields on $K$. Following this work, investigating the homotopy type of $\mathcal{M}(K)$ has been a major task in discrete Morse theory [1, 5, 6, 15].
This paper continues the study and clarifies some ambiguities in [17] by following the idea of birth-death transitions to connect discrete Morse functions on $\mathcal{M}(K)$. Precisely, we show in Theorem 4.6 that any two discrete Morse functions $f_{1}$ and $f_{2}$ can be transformed into one another with a finite sequence of birth-death transitions. Moreover, we apply this theorem to give an alternative, self-contained proof of some Forman’s fundamental results in Theorem 5.1 and Theorem 5.2.
This paper is organized as follows. In Section 2, we provide a concise overview of the preliminaries, including the fundamentals of discrete Morse theory, as well as the concepts of connectedness homomorphisms. In Section 3, we introduce the birth-death transitions introduced in [17] between two discrete Morse complexes. In Section 4, we prove Theorem 4.6. Section 5 applies this framework to re-establish key theorems of Forman through the compositions of birth-death transitions. Also, we provide some remarks about the topology of the space of discrete Morse functions on $K$.
2. Preliminaries
This section provides the foundational concepts necessary for the remainder of the paper.
2.1. Discrete Morse Functions and Gradient Vector Fields
For a comprehensive understanding of discrete Morse theory and discrete Morse homology, readers are referred to [7, 8, 12].
Let $K$ be a simplicial complex, $K_{p}$ be the set of $p$-dimensional simplices. We write $\sigma^{(p)}$ to represent a $p$-dimensional simplex, omitting the superscript $p$ when the dimension is clear from context. We write $\tau\prec\sigma$ and $\tau\succ\sigma$ if $\tau$ is a face of $\sigma$, and $\sigma$ is a face of $\tau$.
A discrete Morse function $f:K\to\mathbb{R}$ satisfies for every simplex $\sigma\in K_{p}$,
  1. (1)
    $\#\{\tau^{(p+1)}\succ\sigma\,|\,f(\tau)\leq f(\sigma)\}=0\text{ or }1;$
  2. (2)
    $\#\{v^{(p-1)}\prec\sigma\,|\,f(\sigma)\leq f(v)\}=0\text{ or }1,$
where #\# denotes the cardinality of a set.
A simplex $\sigma^{(p)}$ is called a critical simplex with respect to $f$ if:
  1. (1)
    $\#\{\tau^{(p+1)}\succ\sigma\,|\,f(\tau)\leq f(\sigma)\}=0;$
  2. (2)
    $\#\{v^{(p-1)}\prec\sigma\,|\,f(\sigma)\leq f(v)\}=0.$
If $\sigma$ is critical, then we call $f(\sigma)$ a critical value of $f$. Conversely, if a simplex $\sigma$ is not critical, then we say that $\sigma$ is a non-critical simplex, and $f(\sigma)$ is a non-critical value. The set of $f$-critical $q$-simplices is denoted $Cr_{q}(f):=\{\sigma\in K_{q}\,|\,\sigma^{(q)}\text{ is }f\text{-critical}\}$ .
The gradient vector field $V$ corresponding to $f$ is the collection of pairs $(\alpha^{(p)},\beta^{(p+1)})$ satisfying $\alpha\prec\beta$ and $f(\alpha)\geq f(\beta)$. We call such a pair $(\alpha,\beta)$ a gradient pair.
A gradient $V$-path is a sequence of simplices
$\alpha_{0}^{(p)},\beta_{0}^{(p+1)},\alpha_{1}^{(p)},\beta_{1}^{(p+1)},\cdots,\beta_{r}^{(p+1)},\alpha_{r+1}^{(p)}$
such that for each $i=0,\cdots,r$:
  • $(\alpha^{(p)},\beta^{(p+1)})\in V$,
  • $\beta_{i}^{(p+1)}\succ\alpha_{i+1}^{(p)}$ and $\alpha_{i+1}^{(p)}\neq\alpha_{i}^{(p)}.$
We denote a gradient $V$-path $\gamma$ from $\alpha_{0}^{(p)}$ to $\alpha_{r+1}^{(p)}$, consisting of $p$- and $(p+1)$-dimensional simplices by
[…]
Alternatively, a $V$-path may begin at a $(p+1)$-dimensional simplex $\beta_{0}$ and end at $\beta_{r}$, denoted:
[…]
2.2. Discrete Morse Complexes
As a discrete analog of the Morse-Smale complex in smooth Morse theory (see [14, 9]), the discrete Morse complex is defined by Forman [7]. Its chain groups are generated by critical simplices, with boundaries determined by counting gradient paths weighted by incidence numbers.
Assume each simplex of $K$ is oriented. For $\alpha^{(q-1)}\prec\sigma^{(q)}$, we use $\sigma:\alpha]=\pm 1$ to represent the incidence number of $\sigma$ and $\alpha$, where the sign is chosen according to the orientation.
Let $f:K\to\mathbb{R}$ be a discrete Morse function. The $q$-th chain group of discrete Morse complex $C_{q}^{f}=C_{q}^{f}(K)$ is the free Abelian group generated by $\sigma_{i}\in Cr_{q}(f)$, The $q$-th boundary homomorphism $\partial_{q}^{f}:C_{q}^{f}\to C_{q-1}^{f}$ is defined as the linear extension of
$\partial_{q}^{f}(\sigma_{i}):=\sum_{\alpha\in Cr_{q-1}(f)}n^{f}(\sigma_{i},\alpha)\alpha.$
Here, $n^{f}(\sigma_{i},\alpha)$ is given by
$n^{f}(\sigma_{i},\alpha):=\sum_{\gamma}m(\gamma),$
[…]
It is known that $\partial^{f}\circ\partial^{f}=0$ [7, Theorem 7.1]. Hence, $\{C_{\ast}^{f},\partial_{\ast}^{f}\}$ forms the discrete Morse chain complex, and the discrete Morse homology is defined as
$H_{\ast}^{f}:=\frac{\text{ker}\partial^{f}_{\ast}}{\text{im}\partial^{f}_{\ast}}.$
Moreover, discrete Morse homology is isomorphic to the ordinary homology [7, Theorem 7.1].
2.3. Connectedness and Connectedness Homomorphisms
Let $f_{1},f_{2}:K\to\mathbb{R}$ be distinct discrete Morse functions on $K$, with respective gradient vector fields $V_{1}$ and $V_{2}$. The notion of connectedness between critical simplices $\tilde{\sigma}_{1}\in Cr_{q}(f_{1})$ and $\tilde{\sigma}_{2}\in Cr_{q}(f_{2})$ is first introduced by King, Mramor Kosta, and Knudson [10]. They applied the concept to investigate the structure of the gradient vector fields and provided some interesting applications.
Precisely, for critical simplices $\tilde{\sigma}_{1}\in Cr_{q}(f_{1})$ and $\tilde{\sigma}_{2}\in Cr_{q}(f_{2})$, we say that $\tilde{\sigma}_{1}$ is connected to $\tilde{\sigma}_{2}$ if
[…]
Note that the corresponding connectedness coefficient $n^{f_{1}}(\tilde{\sigma}_{1},\tilde{\sigma}_{2})$ captures the number (weighted by incidence) of such gradient paths.
Furthermore, we say that $\tilde{\sigma}_{1}$ is strongly connected to $\tilde{\sigma}_{2}$ if $\tilde{\sigma}_{1}$ is connected to $\tilde{\sigma}_{2}$ and $\tilde{\sigma}_{2}$ is connected to $\tilde{\sigma}_{1}$, denoted $\tilde{\sigma}_{1}\leftrightarrow\tilde{\sigma}_{2}.$
For each $q\geq 1$, the $q$-dimensional connectedness homomorphism from $C_{q}^{f_{1}}$ to $C_{q}^{f_{2}}$
$h_{q}:C_{q}^{f_{1}}\to C_{q}^{f_{2}}$
is defined as the linear extension of
$h_{q}(\tilde{\sigma}_{1}):=\sum_{\sigma_{j}\in Cr_{q}(f_{2})}n^{f_{1}}(\tilde{\sigma}_{1},\sigma_{j})\sigma_{j},$
where $\tilde{\sigma}_{1}\in Cr_{q}(f_{1})$.
For $q=0$, inheriting the 0-dimensional connectedness between simplices $\tilde{\sigma}_{1}^{(0)}$ and $\tilde{\sigma}_{2}^{(0)}$, we let
$h_{0}:C_{0}^{f_{1}}\to C_{0}^{f_{2}}$
be the linear extension of
$h_{0}(\tilde{\sigma}_{1}):=\sum_{\sigma_{j}\in Cr_{0}(f_{2})}n^{f_{2}}(\tilde{\sigma}_{1},\sigma_{j})\sigma_{j},$
where $\tilde{\sigma}_{1}\in Cr_{0}(f_{1})$.
On the other hand, we can also define the reverse connectedness homomorphism from $C_{q}^{f_{2}}$ to $C_{q}^{f_{1}}$ with a parallel way as
$g_{q}:C_{q}^{f_{2}}\to C_{q}^{f_{1}}.$
Note, however, that these maps are generally not chain maps, since connectedness is not transitive in general.
3. Birth-death transitions between discrete Morse complexes
In this section, we review and provide more details on birth-death transitions introduced in [17]. The notion of birth-death transitions is a discrete analogue inspired by classical Morse theory’s creation and cancellation of critical points. These transitions act as chain maps between discrete Morse complexes $C_{\ast}^{f_{1}}$ and $C_{\ast}^{f_{2}}$ adding or removing a pair of critical simplices in consecutive dimensions, while preserving the underlying homology [17]. Building upon the connectedness homomorphisms defined previously, we now formalize the birth and death transitions and investigate their structural properties.
Fix a dimension $i$. Suppose two discrete Morse functions $f_{1},f_{2}:K\to\mathbb{R}$ on a simplicial complex $K$ such that:
$Cr_{i}(f_{1})=\#Cr_{i}(f_{2})-1;\quad\#Cr_{i-1}(f_{1})=\#Cr_{i-1}(f_{2})-1,$
and for all $q\neq i,i-1$,
$Cr_{q}(f_{1})=Cr_{q}(f_{2}).$
We denote the two additional $f_{2}$-critical simplices simplices as $\tilde{\sigma}\in Cr_{i}(f_{2})$ and $\tilde{\alpha}\in Cr_{i-1}(f_{2})$, with the coefficient $n^{f_{2}}(\tilde{\sigma},\tilde{\alpha})=k\in\mathbb{Z}.$
Definition 3.1.
Under the conditions specified above, we define
  • The connectedness homomorphism $h:C_{q}^{f_{1}}\to C_{q}^{f_{2}}$ as a birth transition;
  • The connectedness homomorphism $g:C_{q}^{f_{2}}\to C_{q}^{f_{1}}$ as a death transition.
In [17], the following properties were assumed for simplicity. Here we state them explicitly as lemmas to clarify their behaviours. Note that these lemmas simply follow from the assumptions in the definition of birth-death transitions.
Lemma 3.2.
Let $h:C_{q}^{f_{1}}\to C_{q}^{f_{2}}$ and $g:C_{q}^{f_{2}}\to C_{q}^{f_{1}}$ be a birth transition and a death transition, respectively.
Then,
  • For any $\delta_{1}\in Cr_{q}(f_{1})$,
    $h_{q}(\delta_{1})=\begin{cases}\delta_{1}&\text{ if }q\neq i;\\ \delta_{1}+n^{f_{1}}(\delta_{1},\tilde{\sigma})\tilde{\sigma}&\text{ if }q=i.\par\par\end{cases}$
  • For any $\delta_{2}\in Cr_{q}(f_{2})$,
[…]
Lemma 3.3.
Let $h:C_{q}^{f_{1}}\to C_{q}^{f_{2}}$ and $g:C_{q}^{f_{2}}\to C_{q}^{f_{1}}$ be a birth transition and a death transition, respectively.
Then, with respect to the boundary homomorphisms $\partial_{\ast}^{f_{1}}$ and $\partial_{\ast}^{f_{2}}$, the following equations (1) to (5) hold.
Namely, when $q\neq i,i+1\text{ and }\delta\neq\tilde{\alpha}$,
(1) $\partial^{f_{1}}_{q}(\delta)=\partial^{f_{2}}_{q}(\delta).$
For any $\tau\in Cr_{i+1}(f_{2})$,
(2) $\partial_{i+1}^{f_{2}}(\tau)=\partial_{i+1}^{f_{1}}(\tau)+\sum_{\sigma\in Cr_{i}(f_{1})}n^{f_{1}}(\tau,\sigma)\cdot n^{f_{1}}(\sigma,\tilde{\sigma})\tilde{\sigma}$.
For any $\sigma\in Cr_{i}(f_{2})$ and $\sigma\neq\tilde{\sigma}$,
(3) $\partial_{i}^{f_{2}}(\sigma)=\partial_{i}^{f_{1}}(\sigma)+n^{f_{1}}(\sigma,\tilde{\sigma})(k\tilde{\alpha}+\sum_{\alpha\in Cr_{i-1}(f_{1})}k\cdot n^{f_{2}}(\tilde{\alpha},\alpha)\alpha).$
(4) $\partial_{i}^{f_{2}}(\tilde{\sigma})=k\tilde{\alpha}+\sum_{\alpha\in Cr_{i-1}(f_{1})}k\cdot n^{f_{2}}(\tilde{\alpha},\alpha)\alpha.$
(5) $\partial_{i-1}^{f_{2}}(\tilde{\alpha})=\sum_{\alpha\in Cr_{i-1}(f_{2})}n^{f_{2}}(\tilde{\alpha},\alpha)\sum_{\omega\in Cr_{i-2}(f_{1})}n^{f_{1}}(\alpha,\omega)\omega.$
Remark 3.4.
The above lemmas omit the case $q=0$, in which the connectedness coefficient of $n^{f_{1}}$ and $n^{f_{2}}$ may exchange due to the convention regarding 0-dimensional paths (see Section 2.3). However, this does not alter the validity of the results.
Theorem 3.5.
[17, Theorem 4.1]. Let $f_{1},f_{2}:K\to\mathbb{R}$ be discrete Morse functions. Suppose that $h_{q}:C_{q}^{f_{1}}\to C_{q}^{f_{2}}$ and $g_{q}:C_{q}^{f_{2}}\to C_{q}^{f_{1}}$ be a birth transition and a death transition, respectively.
Then, $h$ and $g$ are chain maps.
Besides chain maps, these maps also induce isomorphisms in homology. We show that $h$ and $g$ form a chain-homotopy equivalence as follows.
Theorem 3.6.
Suppose that $h:C_{q}^{f_{1}}\to C_{q}^{f_{2}}$ and $g:C_{q}^{f_{2}}\to C_{q}^{f_{1}}$ be a birth transition and a death transition, respectively.
Then, $h$ and $g$ are chain-homotopy inverses to each other, that is, $g\circ h$ and $h\circ g$ are chain homotopic to the identities.
Proof.
Suppose that two additional $f_{2}$-critical simplices are $\tilde{\sigma}\in\mathrm{Cr}_{i}(f_{2})$ and $\tilde{\alpha}\in\mathrm{Cr}_{i-1}(f_{2})$, with $n^{f_{2}}(\tilde{\sigma},\tilde{\alpha})=k$. For any $f_{1}$-critical simplex $\delta$, one checks immediately that $g\circ h(\delta)=\delta$. Hence
$g\circ h=\mathrm{id}\quad$
on $C^{f_{1}}_{\ast}.$
On $C^{f_{2}}_{\ast}$, the only nontrivial cases are the additional pair $\tilde{\sigma},\tilde{\alpha}$. One computes
$h\circ g(x)=\begin{cases}0,&x=\tilde{\sigma},\\[6.0pt] -\tilde{\alpha}-\sum\limits_{\beta}n^{f_{2}}(\tilde{\alpha},\beta)\,\beta,&x=\tilde{\alpha},\\[10.0pt] x,&\text{otherwise}.\end{cases}$
Define a homotopy
$s:C^{f_{2}}_{\ast}\to C^{f_{2}}_{\ast+1},\qquad s(x)=\begin{cases}-\tfrac{1}{k}\tilde{\sigma},&x=\tilde{\alpha},\\[6.0pt] 0,&\text{otherwise}.\end{cases}$
Then.
$(\partial^{f_{2}}s+s\partial^{f_{2}})(x)=\begin{cases}-\tilde{\sigma},&x=\tilde{\sigma},\\[6.0pt] -\tilde{\alpha}-\sum\limits_{\beta}n^{f_{2}}(\tilde{\alpha},\beta)\,\beta,&x=\tilde{\alpha},\\[10.0pt] 0,&\text{otherwise}.\end{cases}$
Thus
$h\circ g-\mathrm{id}=\partial^{f_{2}}s+s\partial^{f_{2}}.$
Therefore $g\circ h=\mathrm{id}$ and $h\circ g\simeq\mathrm{id}$, so $h$ and $g$ are chain-homotopy inverses. ∎
Lemma 3.7.
Let $f_{1},f_{2}:K\to\mathbb{R}$ be discrete Morse functions and $V_{1},V_{2}$ be the corresponding gradient vector field, respectively. We use $h_{q}:C_{q}^{f_{1}}\to C_{q}^{f_{2}}$ and $g_{q}:C_{q}^{f_{2}}\to C_{q}^{f_{1}}$ to represent the connectedness homomorphisms.
Then, $V_{1}=V_{2}$ if and only if $h=g=\text{id}.$
Proof.
Suppose $V_{1}=V_{2}$. Then, for every critical simplex $\sigma$,
$h_{q}(\sigma)=\sigma=g_{q}(\sigma).$
This follows from the absence of gradient paths between distinct $q$-dimensional critical simplices.
Conversely, if $h=g=\mathrm{id}$, then no new critical simplices are introduced or removed, and the gradient vector fields must coincide. ∎
Example 3.8.
Consider the simplicial complex $K$ with discrete Morse functions $f_{1}$ and $f_{2}$ depicted in Figure 1. We follow the definition to study the connectedness homomorphisms between those two discrete Morse complexes.
[…]
(a) Discrete Morse function $f_{1}$.
[…]
(b) Discrete Morse function $f_{2}$.
Figure 1. A birth-death transition.
Note that the formulation of birth-death transitions and Theorem 3.5 recover, on the chain level, the essential ideas presented in [8, Section 11] regarding the cancellation of critical cells. Our framework provides an explicit algebraic mechanism via birth-death transitions, extending the intuition of smooth Morse theory [13] into the discrete combinatorial setting.
4. Connecting discrete Morse functions via birth-death transitions
In discrete Morse theory, the space of all discrete Morse functions on a simplicial complex $K$ is often studied through their associated gradient vector fields. Two discrete Morse functions $f_{1},f_{2}:K\to\mathbb{R}$ are considered equivalent when their gradient vector fields $V_{1}$ and $V_{2}$ coincide.
The set of all gradient vector fields forms a simplicial complex, known as the complex of discrete Morse functions, denoted $\mathcal{M}(K)$, introduced by Chari and Joswig [4]. This complex parametrizes acyclic matchings (gradient vector fields) in the Hasse diagrams of $K$. In particular, when $K$ is a graph, this simplicial complex coincides with the complex of rooted forests, studied by Kozlov [11].
Also in [4], Chari-Joswig asks how to relate any two acyclic matchings by a sequence of finite steps. We resolve this affirmatively on the chain level, showing that any two discrete Morse functions $f_{1},f_{2}:K\to\mathbb{R}$ can be connected through a finite sequence of birth-death transitions, as defined in Section 3.
We first define the complex of discrete Morse functions.
Definition 4.1.
Given a simplicial complex $K$, we define the complex of discrete Morse functions $\mathcal{M}(K)$ as the collection of subsets of edges in the Hasse diagram of $K$ that form acyclic matchings.
Alternatively, we may define it more explicitly:
We define a simplicial complex $\widetilde{\mathcal{M}}(K)$ on $K$ as follows. We say that a primitive vector field of $K$ is a set $P$ containing one pair of simplices $\alpha$ and $\sigma$ of $K$ such that $\alpha\prec\sigma$ and $\dim\sigma-\dim\alpha=1$.
We let $\mathcal{P}(K)=\{P_{1},P_{2},\ldots,P_{n}\}$ be the set of all primitive vector fields of $K$.
The simplices of $\widetilde{\mathcal{M}}(K)$ are defined as:
  • 0-simplices: $\widetilde{\mathcal{M}}^{0}(K):=\mathcal{P}(K)$;
  • $q$-simplices: $(P_{i_{0}},P_{i_{1}},\ldots,P_{i_{q}})\in\widetilde{\mathcal{M}}^{q}(K)$ if $V=\{P_{i_{0}},P_{i_{1}},\ldots,P_{i_{q}}\}$ form an acyclic matching (i.e., a gradient vector field).
To ensure that all Morse functions are included, we augment $\widetilde{\mathcal{M}}(K)$ with the empty simplex, that is,
$\mathcal{M}(K)=\widetilde{\mathcal{M}}(K)\cup\{\tilde{\emptyset}\}$.
Note that although the empty simplex $\tilde{\emptyset}$ is sometimes ignored in other papers, this inclusion of empty simplex to $\widetilde{\mathcal{M}}(K)$ is necessary in our framework to ensure that $\mathcal{M}(K)$ includes all possible discrete Morse functions, including the discrete Morse functions where all simplices are critical.
Such discrete Morse function always exist; for example: the discrete Morse function
$f:K\to\mathbb{R}$
with
$f(\sigma)=\dim\sigma$
corresponds to the gradient vector field $V_{0}$, in which every simplex is critical.
Recall that in a simplicial complex, the empty simplex $\tilde{\emptyset}$ has dimension 1-1 and is a face of every vertex. It is not hard to verify that if $K$ is a simplicial complex, then $K\cup\{\tilde{\emptyset}\}$ is still an (abstract) simplicial complex. We consider the poset topology of this augmented simplicial complex ordered by inclusion.
As one of the most important statement of Cerf theory, any two Morse functions $\phi_{1},\phi_{2}$ on the same manifold can be related by a finite sequence of moves involving the creation and cancellation of critical pairs. We can use the birth-death transitions to represent a discrete version of this sequence between two discrete Morse functions $f_{1},f_{2}:K\to\mathbb{R}$. We first study the local behavior of such transitions in the complex $\mathcal{M}(K)$.
By $\overline{\sigma}_{1},\overline{\sigma}_{2}\in\widetilde{\mathcal{M}}(K)$, we denote the simplex in $\widetilde{\mathcal{M}}(K)$ corresponding to $V_{1},V_{2}$, respectively.
Lemma 4.2.
If $\overline{\sigma}_{1}\prec\overline{\sigma}_{2}$ and $\dim\overline{\sigma}_{2}-\dim\overline{\sigma}_{1}=1$, then
  • The connectedness homomorphism $h:C^{V_{1}}_{\ast}\to C^{V_{2}}_{\ast}$ is a death transition;
  • The connectedness homomorphism $g:C^{V_{2}}_{\ast}\to C^{V_{1}}_{\ast}$ is a birth transition.
Proof.
Suppose that $\overline{\sigma}_{1}\prec\overline{\sigma}_{2}$ and $\dim\overline{\sigma}_{2}-\dim\overline{\sigma}_{1}=1$. Then, $V_{1}$ differs $V_{2}$ from a primitive vector field $(\alpha,\sigma)$. More precisely, $V_{2}=V_{1}\cup(\alpha,\sigma).$ This implies that both $\alpha$ and $\sigma$ are $V_{1}$-critical.
Note that $(\alpha,\sigma)\in V_{2}$ implies $n^{f_{1}}(\sigma,\alpha)\neq 0$; otherwise, one of the following contradiction arises
  • there is no $V_{1}$-gradient path connecting $\sigma$ and $\alpha$;
  • the coefficients of multiple $V_{1}$-gradient paths cancel with each other.
Case (1) contradicts with either $\alpha\prec\sigma$; and case (2) forms a cycle with no $V_{2}$-critical simplices, which contradicts with the acyclicity of a gradient vector field.
Therefore, by definition, the connectedness homomorphism $g$ is a birth transition, and the connectedness homomorphism $h$ is a death transition. ∎
Example 4.3.
Figure 2 $(b)$ illustrates an example of the complex of discrete Morse functions obtained from the simplicial complex $K$ in $(a)$.
We let $\overline{\sigma}_{1}=(a,e_{1})$ and $\overline{\sigma}_{2}=((a,e_{1}),(b,e_{2}))$. Thus, the corresponding gradient vector fields $V_{1}=\{(a,e_{1})\}$ and $V_{2}=\{(a,e_{1}),(b,e_{2})\}$. Hence, one can check that the connectedness homomorphisms $h$ and $g$ are a death and birth transition, respectively.
[…]
(a) Simplicial complex $K$.
$(a,e_{1})$$(c,e_{2})$$(a,e_{3})$$(b,e_{2})$$(c,e_{3})$$(b,e_{1})$
(b) $\widetilde{\mathcal{M}}(K)$
Figure 2. An example of discrete Morse function complex.
With regard to the empty simplex $\tilde{\emptyset}$, we state the following lemma.
Lemma 4.4.
Let $v$ be any vertex of $\mathcal{M}(K)$, and $V$ be the corresponding gradient vector field.
Then,
  1. (1)
    the connectedness homomorphism $h_{0}:C_{\ast}^{V}\to C_{\ast}^{V_{0}}$ is a birth transition;
  2. (2)
    the connectedness homomorphism $g_{0}:C_{\ast}^{V_{0}}\to C_{\ast}^{V}$ is a death transition,
where $V_{0}$ is the empty gradient vector field.
Proof.
The proof is straightforward from definitions. ∎
We illustrate this in Figure 3 and 4, where the birth-death transitions are drawn in dashed lines.
Example 4.5.
Let $K$ be an edge. $\widetilde{\mathcal{M}}(K)$ consists of two points, thus is not connected. Hence, there is no path of simplices of $\widetilde{\mathcal{M}}(K)$ connecting those two vertices.
By inserting $\tilde{\emptyset}$ to $\widetilde{\mathcal{M}}(K)$, as we do in Fig.4, the path $(a,e_{1})\leftrightarrow\tilde{\emptyset}\leftrightarrow(b,e_{1})$ connects two vertices. And the path represents a composition of a birth transition and a death transition.
[…]
(a) Simplicial complex $K$.
$(a,e_{1})$$(b,e_{1})$
(b) $\widetilde{\mathcal{M}}(K)$
Figure 3. An example of discrete Morse function complex.
[…]
(a) Simplicial complex $K$.
$(a,e_{1})$$(b,e_{1})$$\tilde{\emptyset}$
(b) $\mathcal{M}(K)$
Figure 4. The empty simplex.
Combining the above lemmas, we recover the idea of Cerf theory on $K$ by connecting the discrete Morse functions with birth-death transitions.
Theorem 4.6.
Let $K$ be a finite simplicial complex and $f_{1},f_{2}:K\to\mathbb{R}$ be discrete Morse functions.
Then, $f_{1}$ and $f_{2}$ can be transformed into one another with a finite sequence of birth-death transitions.
Proof.
Let $V_{1}$ and $V_{2}$ be the gradient vector fields corresponding to $f_{1}$ and $f_{2}$. Suppose that these correspond to simplices $\overline{\sigma}_{1},\overline{\sigma}_{2}\in\mathcal{M}(K)$.
Then, we can construct a path consisting of successive adjacent simplices
$\overline{\sigma}_{1}^{(n)}\succ\alpha_{1}^{(n-1)}\succ\alpha_{2}^{(n-2)}\succ\cdots\succ\alpha_{n}^{0}\succ\tilde{\emptyset}$
connecting $\overline{\sigma}_{1}$ to $\tilde{\emptyset}$, denoted $\overline{\sigma}_{1}\;\to\;\tilde{\emptyset}$. Each adjacent pair corresponds to a birth transition as established in Lemma 4.2(2), when $\dim\alpha_{i}\geq 1$; and established in Lemma 4.4(1), when $\dim\alpha_{i}=0$.
Similarly, we can also construct a path connecting $\tilde{\emptyset}$ to $\overline{\sigma}_{2}$, each step corresponding to a death transition.
Hence, both $\overline{\sigma}_{1}$ and $\overline{\sigma}_{2}$ can be connected to $\tilde{\emptyset}$ via such paths:
$\overline{\sigma}_{1}\;\to\;\tilde{\emptyset}\quad\text{and}\quad\tilde{\emptyset}\;\to\;\overline{\sigma}_{2}.$
Concatenating these two paths gives a sequence
$\overline{\sigma}_{1}\;\to\;\tilde{\emptyset}\;\to\;\overline{\sigma}_{2},$
lying entirely in $\mathcal{M}(K)$, in which each adjacent pair of simplices corresponds to a birth or death transition. Since $K$ is finite, this sequence has finitely many steps, completing the proof.
5. An alternative proof of Forman’s Theorems
In Forman’s original theory, a simplicial complex $K$ is simplified through the collapse of gradient pairs (pairs of non-critical simplices) guided by the descending values of the discrete Morse function $f$, or equivalently, the arrows in the associated gradient vector field $V$. This collapsing sequence reduces $K$ to a smaller simplicial complex with critical simplices encoding its homology.
Dually, we can equivalently generate/cancel the critical simplices, the generators of the discrete Morse complex $C_{\ast}^{f}$, in pairs similar to cancellation in classical Morse theory. In this process, $K$ remains unchanged, and the homological structure is modified through algebraic transitions.
Theorem 4.6 establishes that any two gradient vector fields in $\mathcal{M}(K)$ are path-connected via birth-death transitions. Besides the algebraic framework given in algebraic Morse theory, we aim to provide more geometric information and intuitions by the introduction of connectedness homomorphisms and birth-death transitions.
One of the fundamental results in discrete Morse theory, established by Forman [7], is the equivalence between discrete Morse homology and simplicial homology. In our framework, where birth-death transitions act as chain maps between discrete Morse complexes, this correspondence can be recovered naturally.
Theorem 5.1.
[7, Theorem 7.1] Let $K$ be a simplicial complex and $f:K\to\mathbb{R}$ be a discrete Morse function.
Then,
$H_{\ast}(K)\cong H_{\ast}^{\text{simp}}(K),$
where $H_{\ast}^{\text{simp}}(K)$ denotes the simplicial homology of $K$.
Proof.
Let $V_{0}$ be the empty gradient vector field (i.e., the one where all simplices are critical). Define an isomorphism
$\psi_{q}:C_{q}^{\text{simp}}\to C_{q}^{V_{0}},$
for all $q$, where $C_{q}^{\text{simp}}$ denotes the $q$-th simplicial chain group of $K$. Note that
$\psi_{q}=\text{id}_{q},$
for all $q$, as the generators of $C_{q}^{\text{simp}}$ and $C_{q}^{V_{0}}$ coincide.
By Theorem 4.6, the gradient vector field $V$ corresponding to $f$ is connected to $V_{0}$ through a sequence of birth-death transitions, each of which defines a chain homotopy equivalence. Thus, composing the induced maps yields a homology isomorphism:
$H_{\ast}^{V}(K)\cong H_{\ast}^{V_{0}}(K)\cong H_{\ast}^{\text{simp}}(K),$
completing the proof. ∎
Besides Theorem 5.1, other Forman’s theorems can also be recovered in the sense of “collapsing critical simplices”, similar to the proof above. We provide one more example, the Morse equality, as follows.
Theorem 5.2.
[7, Theorem 2.11] Let $K$ be a finite simplicial complex and $f:K\to\mathbb{R}$ be a discrete Morse function.
Then,
$\sum_{q}(-1)^{q}\beta_{q}(K)=\sum_{q}(-1)^{q}Cr_{q}(f),$
where $\beta_{q}(K)$ denotes the $q$-th Betti number of $K$.
Proof.
Let $f_{0}:K\to\mathbb{R}$ be a discrete Morse function whose every simplex is critical. Theorem 4.6 implies that there exists a finite sequence of birth-death transitions transforming $f_{0}$ to $f$. By definition, a birth/death transition generates/cancels a pair of $(q,q+1)$-dimension critical simplices. Thus, the alternating sum remains invariant under such transitions. That is,
$\sum_{q}(-1)^{q}Cr_{q}(f)=\sum_{q}(-1)^{q}Cr_{q}(f_{0}).$
Moreover, since
$\sum_{q}(-1)^{q}Cr_{q}(f_{0})=\chi(K)=\sum_{q}(-1)^{q}\beta_{q},$
where $\chi(K)$ denotes the Euler characteristic of $K$, we have the desired equality.
Several remarks on topology of the set of discrete Morse functions.
Recall that in [17], we introduced the following equivalence relation on the set of all discrete Morse functions on a simplicial complex $K$, denoted, $\mathcal{F}$, to study its topological properties.
Definition 5.3.
Let $f_{1},f_{2},\ldots,f_{n}:K\to\mathbb{R}$ be a sequence of discrete Morse functions, with corresponding discrete Morse complexes $C_{\ast}^{f_{1}},C_{\ast}^{f_{2}},\ldots,C_{\ast}^{f_{n}}$. If, for each $i=1,2,\cdots,n-1$, the connectedness homomorphism $h^{i}_{\ast}:C_{\ast}^{f_{i}}\to C_{\ast}^{f_{i+1}}$ is either an isomorphism or a birth/death transition, then we call the sequence of discrete Morse functions a birth-death transition sequence. Any two discrete Morse functions $f_{j}$ and $f_{k}$ in such a sequence are called transitively connected, denoted $f_{j}\stackrel{{\scriptstyle T}}{{\sim}}f_{k}$.
This defines a natural equivalence relation:
Proposition 5.4.
Transitive connectedness $\stackrel{{\scriptstyle T}}{{\sim}}$ is an equivalence relation on $\mathcal{F}$.
Proof.
Reflexivity holds since $C_{\ast}^{f}\cong C_{\ast}^{f}$ via the identity map. Symmetry follows from the reversibility of birth-death transitions and the invertibility of isomorphisms. Transitivity is built into the definition of the sequence.
This equivalence relation allows us to characterize the topology of $\mathcal{F}$ via the quotient $\mathcal{F}/\stackrel{{\scriptstyle T}}{{\sim}}$.
Theorem 5.5.
Let $K$ be a simplicial complex and $\mathcal{F}$ be the set of all discrete Morse functions on $K$.
Then,
$\mathcal{F}/\stackrel{{\scriptstyle T}}{{\sim}}=\{\ast\}.$
Proof.
By Theorem 4.6, any two discrete Morse functions $f_{1},f_{2}\in\mathcal{F}$ can be connected by a finite sequence of birth-death transitions. Each transition $h_{\ast}^{i}:C_{\ast}^{f_{i}}\to C_{\ast}^{f_{i+1}}$ is chain map, and isomorphisms arise when gradient vector fields are identical (Lemma 3.7). Thus, $f_{1}\stackrel{{\scriptstyle T}}{{\sim}}f_{2}$ for all $f_{1},f_{2}\in\mathcal{F}$, implying the quotient consists of a single point. That is,
$\mathcal{F}/\stackrel{{\scriptstyle T}}{{\sim}}=\{\ast\}.$
Corollary 5.6.
Let $K$ be a simplicial complex. Then, the augmented complex of discrete Morse functions $\mathcal{M}(K)$ is connected with poset topology.
Remark 5.7.
Note that higher-dimensional homology of $\mathcal{M}(K)$ and $\widetilde{\mathcal{M}}(K)$ are isomorphic, as $\tilde{\emptyset}$ affects only connectivity at dimension 0. This viewpoint aligns with the results of Scoville and Zaremsky [15, Theorem 5.4]. They proved that almost all $\widetilde{\mathcal{M}}(K)$ (except Example 4.5) are connected, which implies Corollary 5.6.